The risk-free rate is 5.3 percent and the market risk premium is 5 percent. Assume that required returns are based on the CAPM. Your $1 million portfolio consists of $ 568 ,000 invested in a stock that has a beta of 1.8 and the remainder invested in a stock that has a beta of 1.5 . What is the required return on this portfolio? Enter your answer to the nearest .1%. Do not use the % sign in your answer, thus 12.1% is 12. 1 rather than 12.1% or .121.
Answer #1

Investment in beta of 1.5 stock=(1,000,000-568000)=$432000

Portfolio beta=Respective beta*Respective investment weight

=(568000/1,000,000*1.8)+(432000/1,000,000*1.5)

=1.6704

Hence required return of portfolio=risk free rate+beta*market risk premium

=5.3+(1.6704*5)

=13.7(Approx).